OpenStudy (anonymous):

Can someone solve this Equation 9+n/5=19

OpenStudy (lgbasallote):

cross multiply first 9 +n = 95 n = 86

OpenStudy (lgbasallote):

^that's assuming that it's (9+n)/5 = 19. However, if 5 is only under n then that would be different

OpenStudy (anonymous):

9 + n/5 = 19 Subtract 9 from both sides n/5 = 10 Multiply both sides by 5, to isolate n n = 50.